explain the chemical reaction of charging lead acid battery

Grade:10

1 Answers

Umakant biswal
7 years ago
it works by reverse chemical reaction to store energy , actually they contain lead plates dipped in dillute sulphuric acid solution and the reaction changes the potential energy into electrical energy . 
its a secondary battery and we can recharge it for 2nd time use also . 
at anode 
pb + so4 (2-) = pbso4 + 2 electron 
at cathode 
pbo2 + so4 (2-) 4 h+ + 2 electron = pbso4 + 2 h20 
HOPE IT CLEARS 
ALL THE BEST ..